OpenStudy (amy0799):

f(x)=(x+8)^2+3 Find the point on the graph of function that is closest to the point (-2, 6).

OpenStudy (danjs):

both the given point, and th epoint on the function would contain the same line, Normal to the curve right

OpenStudy (baru):

i think you will have to write another function for distance distance between f(x) and given point \[\sqrt{(x+2)^2+(f(x)-6)^2}\]

OpenStudy (danjs):

yes, the direction should be normal to the curve,

OpenStudy (baru):

we can square the expression, because min will occur when the expression under the square root is minimum

OpenStudy (danjs):

yeah minimize the distance

OpenStudy (baru):

\[d={(x+2)^2+(f(x)-6)^2}\]

OpenStudy (baru):

just set the derivative of that=0

OpenStudy (danjs):

they give you the function y = place tha ty value in to the distance function , to get distance in terms of only x

OpenStudy (danjs):

if you doin algebra

OpenStudy (amy0799):

f\[f(x)=x^4+32x^3+379x^2+1956x+3725\]

OpenStudy (amy0799):

Is that correct?

OpenStudy (baru):

yea, i guess, dont call it f(x) you might confuse yourself, call it s(x) or something else

OpenStudy (danjs):

\[d = \sqrt{(x+2)^2+(f(x)-6)^2}\] from above replace f(x), for the function value, f(x) =(x+8)2 + 3

OpenStudy (danjs):

simplify all that, so you have distance d in terms of just x,

OpenStudy (danjs):

minimize that function,

OpenStudy (baru):

you can minimiz d^2 instead of d \[d^2=x^4+32x^3+379x^2+1956x+3725\] is what amy is getting

OpenStudy (baru):

differentiate, equate to zero, solve and find x, and corresponding f(x)

OpenStudy (baru):

you might need a calculator :P

OpenStudy (amy0799):

I want to make sure the function i got is correct first, is it?

OpenStudy (baru):

or @DanJS has got me thinking, at the minimum point, the tangent to the curve will be perpendicular to the line connecting point on curve and (-2,6)

OpenStudy (baru):

yea, ur function looks correct

OpenStudy (danjs):

yeah , use the tangent at that point (x,y) and the normal line will have slope -1/y'

OpenStudy (danjs):

so just take the derivative, f'(x), that is the tangent slope at the point (x,y) the slope of the line between (-2,6) and (x,y) on the function, has a slope perpendicular to the tangent slope normal to curve slope = -1/y'

OpenStudy (amy0799):

\[s'(x)=-1/4x^3+96x^2+758x+1956\]

OpenStudy (danjs):

by the way , i just put that distance function d(x) into the calc, and found the d ' (x), and solved that for being 0, x=6

OpenStudy (danjs):

back to the derivative thing slope of normal line is -1/y ' or -1/(2x + 16 and the components of that slope in x and y directions, forming the right triangle between the points are y-6 and x+2 the slope of that normal -1/(2x+16) = (y - 6) / x + 2)

OpenStudy (baru):

How did u get -1/4

OpenStudy (amy0799):

how did you get 2x+16?

OpenStudy (danjs):

the derivative of the function y ' = 2(x+8)^1*1 = 2x + 16

OpenStudy (amy0799):

i thought you're taking the derivative of (x+2)^2+((x+8)^2+3-6)^2

OpenStudy (baru):

\[s(x)=x^4+32x^3+379x^2+1956x+3725\] just take the derivative of this; other stuff was just an alternate method

OpenStudy (danjs):

that is the slope of the tangent at a point on the function the slope of the line between that point (x,y) and (-2,6) will be perpendicular to that tangent line -1/y ' or -1/(2x+16)

OpenStudy (baru):

\[s'(x)=-1/4x^3+96x^2+758x+1956\] this is what you got right?

OpenStudy (baru):

how did you get the -1/4 at the start?

OpenStudy (danjs):

that should be the same as change in y over change in x. or (y - 6 ) / (x + 2)

OpenStudy (danjs):

\[\large \frac{ -1 }{ 2x+16 } = \frac{ y - 6 }{ x + 2 }\] replace the function value for the y variable

OpenStudy (danjs):

\[\large \frac{ -1 }{ 2x+16 } = \frac{ [(x+8)^2 + 3] - 6 }{ x + 2 }\] that solves to x = -6 again, same as the distance formula wau

OpenStudy (danjs):

point on the curve (x,y) = (-6 , f(-6))

OpenStudy (amy0799):

so plug -6 into the original function?

OpenStudy (amy0799):

i got the answer, thank you for the help

OpenStudy (danjs):

the derivative of that distance, is a cubic function right?, that is hard to solve , more than just doing -1/y'(x) = (y - y1) / (x - x1)

OpenStudy (amy0799):

I understand now. (-6,7)